Firm A has price elasticity of demand of –1.5 and a marginal cost of $30. Firm...

  1. Firm A has price elasticity of demand of –1.5 and a marginal cost of $30. Firm B has a price elasticity of demand of –2.0 and a marginal cost of $30. What is the profit maximizing price of each firm?
Answer #1

The markup rule can be used to determine the profit-maximizing price given the marginal cost and the price elasticity of demand. The formula is given below:

1 + ed y MC psed

Where P is the profit-maximizing price, e​​​​​​d​​​ is the price elasticity of demand and MC is the marginal cost.

Given:

Firm A:

Price elasticity of demand = -1.5

Marginal Cost (MC) = $30

Firm B:

Price elasticity of demand = -2

Marginal Cost = $30

The profit-maximizing price for both the firms is calculated below:

Firm A:

P-1+(-1.5) . -1.5% x 30 = -0.5 - 15 X 30 = 10

Firm B:

p_1+-2) x 30 = =+ x 30 = 15

Therefore, the profit-maximizing price of firm A is $10 and the profit-maximizing price of firm B is $15.